OpenStudy (anonymous):

Write 121^1/2=11 as a logarithmic function.

OpenStudy (anonymous):

121*1/2=11

OpenStudy (dtan5457):

You need to know your base, answer, and exponent.

OpenStudy (dtan5457):

Base=121 exponent=1/2 answer=11

OpenStudy (anonymous):

Thank you guys

OpenStudy (dtan5457):

not finished lol

OpenStudy (dtan5457):

\[1/2=\log_{121}11\]